英語 (English)

発信型英語 (Communicative English) (工 ((化 B ー 1)2 年))

フロストダン · 非常勤講師 2 単位後期火 1 ・ 2

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】 Let the students recognize the global value and necessity of English by presenting topics relevant to modern life. In this way expand the student’s grasp of spoken English and their ability to respond appropriately.

【授業の概要】 The course will consist of a series of exercises to enforce the stated goal. These exercises will be derived from the textbook as well as from outside materials.

【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】 Through the use of ESL activities build up the skill and confidence

level of the students 【授業の計画】

1. Class weekly-schedule: We will follow a systematic overview of important spoken patterns.Speed depends on class comprehension and level. Generally we strive to cover one pattern over 2 classes but may repeat if needed. The patterns will work in a cumulative manner. This schedule will also depend on collateral ESL activities undertaken.

2. Coursework: Assignments will be given based on the need for reinforcement.These regular tasks count toward the final grade. Optional assignments may be added or suggested.

3. An outline of patterns: Self Introduction: present-past. Travel: Past Perfect. Transportation and Services: Adverbs, Wh questions. Travel: Future and necessity. Instructions: Sequences. Contrasts: Past Present Comparisons. Jobs, Machines, and Descriptions; Things we do and have.

【教科書】 Class handouts and other materials to be given at the appropriate points. 【参考書等】 [参考資料] 【成績評価の方法】 Evaluation is based on attendance, participation in class, completion

of homework, and examination 【再試験の有無】 Consideration will be given for reexamination. Meritorious work will also be accepted in lieu of missed assignments and will affect the final grade.

【受講者へのメッセージ】 Please bring dictionary, notebook, pen,pencil, and red pencil. Pocket

translators may also be used.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221220 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ フロスト .

English

Communicative English (工 ((化 B ー 1)2 年)) Dan Frost · Part-time Lecturer 2 units 後期火 1 ・ 2

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ Let the students recognize the global value and necessity of English by presenting topics relevant to modern life. In this way expand the student’s grasp of spoken English and their ability to respond appropriately.

Outline⟩ The course will consist of a series of exercises to enforce the stated goal. These exercises will be derived from the textbook as well as from outside materials.

�������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ Through the use of ESL activities build up the skill and confidence level of

the students

Schedule⟩

1. Class weekly-schedule: We will follow a systematic overview of important spoken patterns.Speed depends on class comprehension and level. Generally we strive to cover one pattern over 2 classes but may repeat if needed. The patterns will work in a cumulative manner. This schedule will also depend on collateral ESL activities undertaken.

2. Coursework: Assignments will be given based on the need for reinforcement.These regular tasks count toward the final grade. Optional assignments may be added or suggested.

3. An outline of patterns: Self Introduction: present-past. Travel: Past Perfect. Transportation and Services: Adverbs, Wh questions. Travel: Future and necessity. Instructions: Sequences. Contrasts: Past Present Comparisons. Jobs, Machines, and Descriptions; Things we do and have.

Textbook⟩ Class handouts and other materials to be given at the appropriate

points. Reference⟩ [参考資料] Evaluation Criteria⟩ Evaluation is based on attendance, participation in class,

completion of homework, and examination Re-evaluation⟩ Consideration will be given for reexamination. Meritorious work will also be accepted in lieu of missed assignments and will affect the final grade.

Message⟩ Please bring dictionary, notebook, pen,pencil, and red pencil. Pocket

translators may also be used.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221220 Contact (Office-Hour, Room, E-mail)⟩

⇒ Frost .